After riding at a steady speed for 40 miles, a bicyclist had a flat tire and walked 5 miles to a repair shop. The cyclist rate was 4 times faster than the walking rate. If the time spend on the cycling and walking was 5 hours, at what rate was the cyclist riding?
Answer by mananth(12270)

After riding at a steady speed for 40 miles, a bicyclist had a flat tire and walked 5 miles to a repair shop. The cyclist rate was 4 times faster than the walking rate. If the time spend on the cycling and walking was 5 hours, at what rate was the cyclist riding?
let walking rate x
cycling rate will be 4x
time taken to ride 40 miles = 40/4x
time taken to walk 5 miles = 5/x
the total time taken was 5 hours
40/4x + 5/x = 5
40+20 =20x
60=20x
x=3 mph the walking rate
riding rate = 4x= 12 mph